OpenStudy (bananas1234):

Which number is halfway between 1/3 and 1/2? 2/12 3/12 5/12 7/12

OpenStudy (mathstudent55):

Take the average of the numbers.

OpenStudy (bananas1234):

What numbers?

OpenStudy (triciaal):

use equivalent fractions then place on the number line 1/3 = x/12 1/2 = y/12 (x + y)/2 where x and y are whole numbers

OpenStudy (mathstudent55):

The only two numbers you are given.

OpenStudy (mathstudent55):

Not the choices.

OpenStudy (bananas1234):

is 5/12 correct? @triciaal

OpenStudy (triciaal):

yes

OpenStudy (bananas1234):

Thank you!

OpenStudy (triciaal):

you are welcome

OpenStudy (bananas1234):

:)